# suppReferences

### v2

@@ Abstract
Real-world misinformation can be partially correct and even factual but misleading. It undermines public trust in science and democracy, particularly on social media, where it can spread rapidly. High-quality and timely correction of misinformation that identifies and explains its (in)accuracies has been shown to effectively reduce false beliefs. Despite the wide acceptance of manual correction, this is difficult promptly correct newly created misinformation and to scale this approach, a concern as technologies like large language models (LLMs) make misinformation easier to produce. LLMs also have versatile capabilities that could accelerate misinformation correction—however, they struggle due to a lack of recent information, a tendency to produce false content, and limitations in addressing multimodal information. We propose Muse, an LLM augmented with access to and credibility evaluation of up-to-date information. By retrieving evidence as refutation or supporting context, Muse identifies and explains (in)accuracies in a piece of content—not presupposed to be misinformation—with references. It also describes images and conducts multimodal searches to verify and correct multimodal content. Fact-checking experts evaluate responses to social media content that are not presupposed to be (non-)misinformation but broadly include incorrect, partially correct, and correct posts, that may or may not be misleading. We propose and evaluate 13 dimensions of misinformation correction quality, ranging from the accuracy of identifications and factuality of explanations to the relevance and credibility of references. The results demonstrate Muse’s ability to promptly write high-quality responses to potential misinformation on social media—overall, Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29%. This work reveals LLMs’ potential to help combat real-world misinformation effectively and efficiently.
@@ Introduction
Muse’s evaluation was based on X Community Notes data. Community Notes empowers people on X, often laypeople, to collaboratively fact-check tweets, which has shown to reduce the spread of misinformation[17]. Every laypeople’s free-response fact-check is associated with a helpfulness score by aggregating the assessments of people with diverse backgrounds, e.g., different political ideologies. A response with a sufficiently high helpfulness score is then displayed on the corresponding tweet and publicly visible[17] (Supplementary Fig. S6). We included the tweets from Community Notes with at least two responses (#=247); one has a high helpfulness score and the other has an average helpfulness score (as of February 2023; Methods). Though we do not presuppose the accuracy of the tweets in Muse’s design and evaluation, we found that more than half of the tweets are not fully (in)accurate or misleading but frequently presented in a way that combines accurate claims and inaccurate or misleading claims (Methods). We further generated responses to these tweets based on Muse and GPT-4 (as of June 2023) (Discussion; Methods). Experts in fact-checking and journalism evaluated the quality of responses by various approaches to the same tweet (Methods); they were blinded to which approach had generated each response. The evaluation contains 13 specific criteria, covering how well a response identifies and explains (in)accuracies, the quality of generated text, and the quality of references (Methods). It also contains the overall quality of a response by taking all 13 evaluation criteria into account (Methods).
@@
Examining specific components of response quality, results show that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in identifying and explaining (in)accuracies (Fig. 2b-f). Experts assessed that Muse-generated responses more explicitly identify and explain where and why a tweet is (in)accurate than GPT-4’s and laypeople’s high-helpfulness responses (Fig. 2b). 89% of Muse’s generated responses explicitly identify and explain (in)accuracies, 16% more than GPT-4-generated responses, 29% more than laypeople’s high-helpfulness responses, and 43% more than laypeople’s average-helpfulness responses (Fig. 2b). As for identifying where a tweet is (in)accurate, we found that Muse more comprehensively identifies a tweet’s (in)accuracies with fewer mistakes—here, mistakes indicate falsely claiming where a tweet should be inaccurate as accurate or where a tweet should be accurate as inaccurate—than GPT-4 and laypeople who produce even high-helpfulness responses (Fig. 2c-d). 91% of Muse’s generated responses have at least one correct identification without any mistake, 11% more than laypeople’s high-helpfulness responses, 19% more than GPT-4-generated responses, and 26% more than laypeople’s average-helpfulness responses (Fig. 2c). Muse has 61% of generated responses accurately identifying all the (in)accuracies in a tweet, vs GPT-4 has 38%, laypeople who produce high-helpfulness responses have 26%, and laypeople who produce average-helpfulness responses have 17% only (Fig. 2d). Furthermore, Muse explains (in)accuracies more precisely and informatively than GPT-4 and laypeople who produce even high-helpfulness responses (Fig. 2e-f). 70% of responses by Muse have fully accurate explanations, vs 55% for laypeople’s high-helpfulness responses, 47% by GPT-4, and 37% for laypeople’s average-helpfulness responses only (Fig. 2e). Meanwhile, the average informativeness score of Muse-generated responses is 8.4, 18% higher than GPT-4-generated responses, 29% higher than laypeople’s high-helpfulness responses, and 50% higher than laypeople’s average-helpfulness responses (Fig. 2f).
@@
Results also demonstrate that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in the quality of generated text (Fig. 2g-k). Muse, when it augments GPT-4 with the capabilities of accessing timely updated knowledge and addressing visuals (Methods), exhibits enhanced relevance ($p=10^{-15}$; $N=460$) and factuality ($p=2\times 10^{-20}$; $N=459$) of text compared to GPT-4 without sacrificing fluency ($p=0.6$; $N=464$), coherence ($p=0.1$; $N=459$), and toxicity ($p=0.8$; $N=464$). Meanwhile, Muse-generated text is more relevant to the responded tweet ($p=2\times 10^{-30}$; $N=464$), factual ($p=4\times 10^{-6}$; $N=464$), fluent ($p=2\times 10^{-10}$; $N=464$), and coherent ($p=10^{-5}$; $N=451$) than the text of high-helpfulness responses by laypeople and additionally less toxic than the text of average-helpfulness responses by laypeople ($p=4\times 10^{-12}$; $N=462$). In particular, Muse-generated text has an average relevance score of 8.7, 18% higher than GPT-4-generated text, 21% higher than the text of high-helpfulness responses by laypeople, and 43% higher than the text of average-helpfulness responses by laypeople (Fig. 2g). 74% of Muse-generated text is completely factual, vs 59% for the text of even high-helpfulness responses by laypeople and 45% for GPT-4-generated text (Fig. 2h). Almost all of Muse-generated text does not have any mistake in the use of English (Fig. 2i) and is not biased, impolite, and provoking (Fig. 2k), and 91% is highly coherent and logical (vs 76% and 61% for laypeople, Fig. 2j).
@@ Discussion
This work also faces the following limitations. First, although Muse is capable of responding to multimodal misinformation with text and images, it cannot accept video inputs. Second, we only focus on English, one of the most spoken languages in the world. Third, we evaluated Muse using real social media content on a single platform, X, as its Community Notes system has shown to reduce the spread of misinformation[17] and transparent. X is also a popular social media platform, and one where more than half of users consume news regularly[46] and where misinformation has been shown to diffuse faster than the truth[13]. Fourth, experts assessed and compared Muse against one other LLM, GPT-4, which can also be seen as an ablation study, since Muse augments GPT-4 (Methods). GPT-4 was chosen as a comparison as it is one of the best performing LLMs across a wide range of tasks available today[47, 48].
@@ Acknowledgements
Table S1
@@
Category
Definition
# of sources
@@
- Very high:
The source is consistently factual, relies on credible information, promptly corrects errors, and has never failed any fact checks in news reporting or opinion pieces.
118
- High:
The source is mostly factual and uses mostly credible, low-biased, or high-factual sources. It corrects errors quickly and has failed only one news fact check and up to two op-ed fact checks.
2,313
- Mostly factual:
The source is generally accurate but may have a few uncorrected fact-check failures. It can fail up to three op-ed fact checks, especially if it is a low-volume site. While it may use biased sources occasionally, it mostly links to factual content. It is usually pro-science but may sometimes use misleading wording or offer alternative viewpoints. It is reasonably transparent and trustworthy most of the time, but caution is advised.
326
- Mixed:
The source may rely on improper sourcing or link to other biased or mixed-factual sources. It often has multiple failed fact checks and does not correct false information or lacks transparency, including the absence of a disclosed mission statement or ownership details. Sources rejecting established scientific consensus on issues such as climate change or vaccines will receive this rating or lower. Sources identified as overt propaganda or designated as hate groups by reputable third-party evaluators will receive this rating or lower due to their inherent bias and potential spread of misleading information.
1,437
- Low:
The source is often unreliable and should be fact-checked for fake news, conspiracy theories, and propaganda.
677
- Very low:
The source is almost always unreliable and should always be fact-checked for intentional misinformation.
252
@@
- Least biased:
The source has minimal bias and uses very few loaded words (i.e., wording that attempts to influence an audience by using an appeal to emotion or stereotypes). It is factual and usually sourced.
1,054
- Left-center:
The source has a slight to moderate liberal bias. It often publishes factual information that utilizes loaded words to favor liberal causes. It is generally trustworthy for information but may require further investigation.
850
- Right-center:
Similar to the definition of left-center bias but replacing liberal with conservative.
492
- (Extremely) left:
The source is moderately to strongly biased toward liberal causes through story selection or political affiliation. It may utilize strong loaded words, publish misleading reports, and omit reporting of information that may damage liberal causes. It may be untrustworthy.
402
- (Extremely) right:
Similar to the definition of (extremely) left bias but replacing liberal with conservative.
314
- Pro-science:
The source consists of legitimate science or is evidence-based through the use of credible scientific sourcing. Legitimate science follows the scientific method, is unbiased, and does not use emotional words. The source also respects the consensus of experts in the given scientific field and strives to publish peer-reviewed science. It may have a slight political bias but adheres to scientific principles.
189
- Conspiracy- pseudoscience:
The source may publish unverifiable information not always supported by evidence. It may be untrustworthy for credible or verifiable information, so fact-checking and further investigation are recommended on a per-article basis when obtaining information from it.
433
- Questionable:
The source exhibits one or more of the following: extreme bias, consistent promotion of propaganda or conspiracies, poor or no sourcing of credible information, a complete lack of transparency, or fake news (i.e., the deliberate attempt to publish hoaxes or disinformation for profit or influence). It may be very untrustworthy and should be fact-checked on a per-article basis.
1,390
- Satire:
The source exclusively uses humor, irony, exaggeration, or ridicule to expose and criticize people’s stupidity or vices, particularly in the context of contemporary politics and other topical issues. It does not attempt to deceive.
148
@@
Figure S28: Impact of post times of tweets on Muse’s performance.

### v3

@@ Abstract
Real-world misinformation can be partially correct and even factual but misleading. It undermines public trust in science and democracy, particularly on social media, where it can spread rapidly. High-quality and timely correction of misinformation that identifies and explains its (in)accuracies has been shown to effectively reduce false beliefs. Despite the wide acceptance of manual correction, it is difficult to be timely and scalable, a concern as technologies like large language models (LLMs) make misinformation easier to produce. LLMs also have versatile capabilities that could accelerate misinformation correction—however, they struggle due to a lack of recent information, a tendency to produce false content, and limitations in addressing multimodal information. We propose Muse, an LLM augmented with access to and credibility evaluation of up-to-date information. By retrieving evidence as refutation or supporting context, Muse identifies and explains (in)accuracies in a piece of content—not presupposed to be misinformation—with references. It also describes images and conducts multimodal searches to verify and correct multimodal content. Fact-checking experts evaluate responses to social media content that are not presupposed to be (non-)misinformation but broadly include incorrect, partially correct, and correct posts that may or may not be misleading. We propose and evaluate 13 dimensions of misinformation correction quality, ranging from the accuracy of identifications and factuality of explanations to the relevance and credibility of references. The results demonstrate Muse’s ability to promptly write high-quality responses to potential misinformation on social media—overall, Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29%. This work reveals LLMs’ potential to help combat real-world misinformation effectively and efficiently.
@@ Introduction
Muse’s evaluation was based on X Community Notes data. Community Notes empowers people on X, often laypeople, to collaboratively fact-check tweets, which has been shown to reduce the spread of misinformation[17]. Every laypeople’s free-response fact-check is associated with a helpfulness score by aggregating the assessments of people with diverse backgrounds, e.g., different political ideologies. A response with a sufficiently high helpfulness score is then displayed on the corresponding tweet and publicly visible[17] (Supplementary Fig. S6). We included the tweets from Community Notes with at least two responses (#=247); one has a high helpfulness score and the other has an average helpfulness score (as of February 2023; Methods). Though we do not presuppose the accuracy of the tweets in Muse’s design and evaluation, we found that more than half of the tweets are not fully (in)accurate or misleading but frequently presented in a way that combines accurate claims and inaccurate or misleading claims (Methods). We further generated responses to these tweets based on Muse and GPT-4 (as of June 2023) (Discussion; Methods). Experts in fact-checking and journalism evaluated the quality of responses by various approaches to the same tweet (Methods); they were blinded to which approach had generated each response. The evaluation contains 13 specific criteria, covering how well a response identifies and explains (in)accuracies, the quality of generated text, and the quality of references (Methods). It also contains the overall quality of a response by taking all 13 evaluation criteria into account (Methods).
@@
Examining specific components of response quality, results show that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in identifying and explaining (in)accuracies (Fig. 2b-f). Experts assessed that Muse-generated responses more explicitly identify and explain where and why a tweet is (in)accurate than GPT-4’s and laypeople’s high-helpfulness responses (Fig. 2b). 89% of Muse’s generated responses explicitly identify and explain (in)accuracies, 16% more than GPT-4-generated responses, 29% more than laypeople’s high-helpfulness responses, and 43% more than laypeople’s average-helpfulness responses (Fig. 2b). As for identifying where a tweet is (in)accurate, we found that Muse more comprehensively identifies a tweet’s (in)accuracies with fewer mistakes—here, mistakes indicate falsely claiming where a tweet should be inaccurate as accurate or where a tweet should be accurate as inaccurate—than GPT-4 and laypeople who produce even high-helpfulness responses (Fig. 2c-d). 91% of Muse’s generated responses have at least one correct identification without any mistake, 11% more than laypeople’s high-helpfulness responses, 19% more than GPT-4-generated responses, and 26% more than laypeople’s average-helpfulness responses (Fig. 2c). Muse has 61% of generated responses accurately identifying all the (in)accuracies in a tweet, vs GPT-4 has 38%, laypeople who produce high-helpfulness responses have 26%, and laypeople who produce average-helpfulness responses have 17% only (Fig. 2d). Furthermore, Muse explains (in)accuracies more precisely and informatively than GPT-4 and laypeople who produce even high-helpfulness responses (Fig. 2e-f). 70% of responses by Muse have fully accurate explanations, vs 55% for laypeople’s high-helpfulness responses, 47% by GPT-4, and 37% for laypeople’s average-helpfulness responses only (Fig. 2e). Meanwhile, the average informativeness score of Muse-generated responses is 7.9, 32% higher than laypeople’s high-helpfulness responses, 36% higher than GPT-4-generated responses, and 65% higher than laypeople’s average-helpfulness responses (Fig. 2f).
@@
Results also demonstrate that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in the quality of generated text (Fig. 2g-k). Muse, when it augments GPT-4 with the capabilities of accessing timely updated knowledge and addressing visuals (Methods), exhibits enhanced relevance ($p=10^{-15}$; $N=460$) and factuality ($p=2\times 10^{-20}$; $N=459$) of text compared to GPT-4 without sacrificing fluency ($p=0.6$; $N=464$), coherence ($p=0.1$; $N=459$), and toxicity ($p=0.8$; $N=464$). Meanwhile, Muse-generated text is more relevant to the responded tweet ($p=2\times 10^{-30}$; $N=464$), factual ($p=4\times 10^{-6}$; $N=463$), fluent ($p=2\times 10^{-10}$; $N=464$), and coherent ($p=10^{-5}$; $N=451$) than the text of high-helpfulness responses by laypeople and additionally less toxic than the text of average-helpfulness responses by laypeople ($p=4\times 10^{-12}$; $N=462$). In particular, Muse-generated text has an average relevance score of 8.7, 18% higher than GPT-4-generated text, 21% higher than the text of high-helpfulness responses by laypeople, and 43% higher than the text of average-helpfulness responses by laypeople (Fig. 2g). 74% of Muse-generated text is completely factual, vs 59% for the text of even high-helpfulness responses by laypeople and 45% for GPT-4-generated text (Fig. 2h). Almost all of Muse-generated text does not have any mistake in the use of English (Fig. 2i) and is not biased, impolite, and provoking (Fig. 2k), and 91% is highly coherent and logical (vs 76% and 61% for laypeople, Fig. 2j).
@@ Discussion
This work also faces the following limitations. First, although Muse is capable of responding to multimodal misinformation with text and images, it cannot accept video inputs. Second, we only focus on English, one of the most spoken languages in the world. Third, we evaluated Muse using real social media content on a single platform, X, as its Community Notes system has been shown to reduce the spread of misinformation[17] and transparent. X is also a popular social media platform, and one where more than half of users consume news regularly[46] and where misinformation has been shown to diffuse faster than the truth[13]. Fourth, experts assessed and compared Muse against one other LLM, GPT-4, which can also be seen as an ablation study, since Muse augments GPT-4 (Methods). GPT-4 was chosen as a comparison as it is one of the best performing LLMs across a wide range of tasks available today[47, 48].
@@ Acknowledgements
Table LABEL:supp:table:mbfc
@@
Category Definition # of sources
@@
- Very high: The source is consistently factual, relies on credible information, promptly corrects errors, and has never failed any fact checks in news reporting or opinion pieces. 118
- High: The source is mostly factual and uses mostly credible, low-biased, or high-factual sources. It corrects errors quickly and has failed only one news fact check and up to two op-ed fact checks. 2,313
- Mostly factual: The source is generally accurate but may have a few uncorrected fact-check failures. It can fail up to three op-ed fact checks, especially if it is a low-volume site. While it may use biased sources occasionally, it mostly links to factual content. It is usually pro-science but may sometimes use misleading wording or offer alternative viewpoints. It is reasonably transparent and trustworthy most of the time, but caution is advised. 326
- Mixed: The source may rely on improper sourcing or link to other biased or mixed-factual sources. It often has multiple failed fact checks and does not correct false information or lacks transparency, including the absence of a disclosed mission statement or ownership details. Sources rejecting established scientific consensus on issues such as climate change or vaccines will receive this rating or lower. Sources identified as overt propaganda or designated as hate groups by reputable third-party evaluators will receive this rating or lower due to their inherent bias and potential spread of misleading information. 1,437
- Low: The source is often unreliable and should be fact-checked for fake news, conspiracy theories, and propaganda. 677
- Very low: The source is almost always unreliable and should always be fact-checked for intentional misinformation. 252
@@
- Least biased: The source has minimal bias and uses very few loaded words (i.e., wording that attempts to influence an audience by using an appeal to emotion or stereotypes). It is factual and usually sourced. 1,054
- Left-center: The source has a slight to moderate liberal bias. It often publishes factual information that utilizes loaded words to favor liberal causes. It is generally trustworthy for information but may require further investigation. 850
- Right-center: Similar to the definition of left-center bias but replacing liberal with conservative. 492
- (Extremely) left: The source is moderately to strongly biased toward liberal causes through story selection or political affiliation. It may utilize strong loaded words, publish misleading reports, and omit reporting of information that may damage liberal causes. It may be untrustworthy. 402
- (Extremely) right: Similar to the definition of (extremely) left bias but replacing liberal with conservative. 314
- Pro-science: The source consists of legitimate science or is evidence-based through the use of credible scientific sourcing. Legitimate science follows the scientific method, is unbiased, and does not use emotional words. The source also respects the consensus of experts in the given scientific field and strives to publish peer-reviewed science. It may have a slight political bias but adheres to scientific principles. 189
- Conspiracy- pseudoscience: The source may publish unverifiable information not always supported by evidence. It may be untrustworthy for credible or verifiable information, so fact-checking and further investigation are recommended on a per-article basis when obtaining information from it. 433
- Questionable: The source exhibits one or more of the following: extreme bias, consistent promotion of propaganda or conspiracies, poor or no sourcing of credible information, a complete lack of transparency, or fake news (i.e., the deliberate attempt to publish hoaxes or disinformation for profit or influence). It may be very untrustworthy and should be fact-checked on a per-article basis. 1,390
- Satire: The source exclusively uses humor, irony, exaggeration, or ridicule to expose and criticize people’s stupidity or vices, particularly in the context of contemporary politics and other topical issues. It does not attempt to deceive. 148
@@
Figure S28: Impact of post times of tweets on the performance of Muse and baselines.